OpenStudy (anonymous):

The perimeter of a parallelogram is 12 ft and its length is 3 ft. The length of the other side is _____ ft.

OpenStudy (ajprincess):

The opposites sides of a parallelogram are equal and parallel.  perimeter = 2(a+b) 12=2(3+b) 12/2=3+b 6=3+b Does that help? @millitarygirl96